the product of two consecutive positive integers is 306. we need to find the integers
Answer by ankor@dixie-net.com(22740)

the product of two consecutive positive integers is 306. we need to find the integers.
:
Let the two integers = x, (x+1)
:
The product:
x(x+1) = 306
:
x^2 + x = 306
:
x^2 + x - 306 = 0
:
An easy way to find the factors, we know the two factors differ by 1
Find the square root of 306 on a calc.
One factor is above this value and one is below.
This would be:
(x+18)(x-17) = 0
x = +17, -18
It wants the positive integers, so our solution will be:
x = 17